# When LLM Judges Inflate Scores: Exploring Overrating in Relevance Assessment

### v2

@@ Abstract.
Human relevance assessment is time-consuming and cognitively intensive, limiting the scalability of Information Retrieval evaluation. This has led to growing interest in using large language models (LLMs) as proxies for human judges. However, it remains an open question whether LLM-based relevance judgments are reliable, stable, and rigorous enough to match humans for relevance assessment. In this work, we conduct a systematic study of overrating behavior in LLM-based relevance judgments across model backbones, evaluation paradigms (pointwise and pairwise), and passage modification strategies. We show that models consistently assign inflated relevance scores — often with high confidence — to passages that do not genuinely satisfy the underlying information need, revealing a system-wide bias rather than random fluctuations in judgment. Furthermore, controlled experiments show that LLM-based relevance judgments can be highly sensitive to passage length and surface-level lexical cues. These results raise concerns about the usage of LLMs as drop-in replacements for human relevance assessors, and highlight the urgent need for careful diagnostic evaluation frameworks when applying LLMs for relevance assessments. Our code and results are publicly available. [^1] https://open-science.anonymous-github.xyz/r/Exploring-Overrating-F11D
@@
$c$s: Information systems Relevance assessment
@@ 1. Introduction and Background
Building reliable relevance annotations for evaluation has long been a central challenge in Information Retrieval (IR). Cranfield-style collections (Cleverdon, 1967) such as TREC DL (Craswell et al., 2020, 2021) rely on trained assessors to annotate query-document pairs, costing substantial time and effort (Soboroff, 2025), while crowdsourced annotations often suffer from limited reliability (Guo et al., 2023; Rahmani et al., 2025; Damessie et al., 2017). Large language models (LLMs) offer a potential alternative by enabling relevance judgments to be generated rapidly and at scale.
Recent studies report that LLM-based relevance judgments can closely align with human assessments. Thomas et al. (2024) employed real searcher feedback to select LLM–prompt configurations aligned with user preferences, demonstrating performance comparable to human assessors in identifying strong systems and difficult queries. These findings were reproduced and extended by Upadhyay et al. (2024a), who further released the open-source UMBRELA framework (Upadhyay et al., 2024b). In subsequent work,  Upadhyay et al. (2025) showed that LLM-generated judgments can replace traditional human labels at the run level in the TREC 2024 RAG Track, with limited benefit from additional human-in-the-loop intervention.
While LLMs have shown to be effective annotation tools, their potential shortcomings as relevance assessors remains insufficiently understood. LLM-based judgments raise several issues. For example, Alaofi et al. (2024) show that LLMs are vulnerable to keyword stuffing, systematically overrating non-relevant passages that contain query terms by assigning inflated relevance labels. Inflated relevance labels — as compared to humans — have been consistently reported in prior work (Upadhyay et al., 2024a; Meng et al., 2025). This behavior also demonstrates how strong aggregate performance can obscure judgment-level errors, which can directly impact downstream evaluation tasks. Moreover, using LLMs as both a retrieval method and as an evaluation (labeling) method introduces circularity, confounding true system improvements with LLM-specific preferences (Clarke and Dietz, 2024). These issues highlight the potential risks when adopting LLMs as relevance assessors.
Since systems are measured over the set of relevance labels for a given set of topics, these labels serve as an upper bound on measurable effectiveness. When relevance judgments are generated by LLMs, this upper bound is implicitly shaped by the models’ own judgment behaviors. LLM-based relevance assessments tend to exhibit label inflation (e.g., assigning a document a higher relevance label than that assigned by a human assessor), which may distort evaluation signals and mask genuine differences in retrieval quality. However, it remains unclear how widespread this behavior is in different model families and evaluation settings.
@@ 2. Overrating Behavior in LLM Judges
Beyond label inflation, we analyze the token-level confidence on the LLM’s predicted label or preference. In the pointwise setting, we determine the prediction of confidently incorrect as a model that assigns near-deterministic probabilities (i.e., softmax probabilities approaching 1.0) to incorrect label tokens, which are measured by the normalized logits on the generation of the label. In the pairwise setting, instability manifests where the LLM yields varying preferences under positional swapping (e.g., $A>B$ vs. $B>A$) while still assigning high confidence to the assessments.
@@ 2.1. Experimental Setup
We conduct experiments on the TREC Deep Learning (DL) Track passage ranking tasks from 2019-2020 (Craswell et al., 2020, 2021), using human relevance labels (0–3) as ground truth. We evaluate four open-weight LLMs: Llama-3.2-3B (Grattafiori and et. al., 2024), Gemma-3-4B (Team, 2025), Mistral-7B (Jiang et al., 2023), and Qwen-3-8B (Yang et al., 2025) under both pointwise and pairwise settings.
@@
Pointwise Evaluation. For pointwise judgments, we measure label inflation through two metrics: (1) Overrate Ratio, the proportion of the instances where the LLM assigns a relevance label exceeding the human label; and (2) Mean Bias, the average signed difference between LLM-assigned labels and human relevance labels (LLM minus Human). We additionally report Cohen’s $\kappa$ to analyze agreement with human judgments. These metrics collectively quantify both the magnitude and distributional patterns of label inflation exhibited by different LLMs. To understand model confidence, we binarized both the human and LLM judgments according to the TREC standard. [^3] Not relevant: 0 and 11; Relevant: $\geq 2$. See: https://trec.nist.gov/data/deep2019.html Then, we aggregate the token-level confidence according to the True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) classes when comparing the human and LLM judgment pairs. Our intuition is that models may be less confident for incorrect (FP or FN) judgments.
Pairwise Evaluation. For pairwise judgment, passages are grouped per query into three categories: Best (top‐graded relevant), Acceptable (relevant, but not the highest grade), and Unacceptable (non‐relevant) (Arabzadeh and Clarke, 2025). We sampled 20 pairs per query from each of the following settings: Best vs. Acceptable; Acceptable vs. Unacceptable; and Best vs. Unacceptable. We treat the higher‐class passage as the ground‐truth winner, and pairs are judged in both presentation orders. We force the LLM to prefer one of the two documents, rather than allowing an explicit tie option. As such, we define a tie as the case where a contradiction arises between the two presentation orderings, meaning that the judgments cannot be linearized.
@@
$2.$ Results: Pointwise Judgment
Model Overrated (%) Mean Bias Cohen’s $\kappa$ Avg Confidence
Bin. Grad. Bin. Grad. Bin. Grad. TP FP TN FN
DL19 Gemma 37.81 61.84 0.341 0.736 0.246 0.107 0.987 0.989 0.989 0.977
Llama 18.82 49.72 0.077 0.564 0.304 0.163 0.902 0.881 0.958 0.914
Mistral 21.53 54.23 0.141 0.557 0.369 0.150 0.939 0.932 0.951 0.904
Qwen 20.62 47.02 0.145 0.459 0.421 0.215 0.977 0.976 0.991 0.986
DL20 Gemma 44.78 66.51 0.431 0.895 0.163 0.088 0.983 0.979 0.986 0.981
Llama 18.57 51.11 0.124 0.691 0.265 0.149 0.908 0.870 0.920 0.871
Mistral 22.32 55.08 0.180 0.646 0.293 0.140 0.941 0.925 0.937 0.910
Qwen 23.42 44.59 0.206 0.516 0.338 0.235 0.977 0.974 0.986 0.968
Table 1. Pointwise overrating metrics on DL2019 and DL2020 under binary and graded relevance labels. Average Confidence is computed under the graded relevance label setting.
Table 1 demonstrates the pervasive nature of overrating across all models and datasets. In the binary setting, all models are substantially more conservative than in the graded setting, yielding markedly lower overrate rates. For instance, Llama’s overrate increases from 18.57% (binary) to 51.11% (graded). This indicates that when restricted to a relevant-irrelevant decision boundary, models apply a stringent relevance criterion. Under the finer-grained graded scale, LLM judges more often promote relevance levels, demonstrating substantial overrating behavior. This shift is also reflected in agreement with human labels: Cohen’s $\kappa$ is consistently lower in the graded setting than in the binary setting.
Examining the label transition matrices in the graded setting reveals that the overrating behavior is dominated by soft errors: non-relevant passages are rarely assigned the highest relevance score, but are frequently misclassified as marginally or moderately relevant. This systematic reluctance to assign the non-relevant label inflates relevance estimates under graded evaluation.
@@ 2.3. Results: Pairwise Judgment
Acc Acc (No Ties) Tie Rate Wrong Wins Corr Incorr
@@
Table 2. Pairwise preference-judgment results on DL2019 and DL2020. Avg Confidence is computed on no-tie cases only: Corr/Incorr are the mean confidence for correct/incorrect preferences.
Table 2 shows the results of the pairwise judgment experiment. Overall, these results reveal universal trends across all evaluated models.
Trend 1: Lack of Discrimination.
As shown in Table 2, the “Tie Rate” is substantial for all models, ranging from $\approx 24\%$ to over 45%. This indicates order-induced inconsistency (preference flips under swapping), not consistent errors where the model repeats the same wrong choice. Furthermore, we investigated how judgment varies with the different graded pairs, as, intuitively, the Best vs Unacceptable pairs should be the easiest to discriminate. However, we observed that all four LLMs tested consistently overrate “Unacceptable” passages, treating them as qualitatively equivalent to the “Best” passages in nearly one-third of cases. Moreover, the tie rate increases systematically as the discrimination task becomes more difficult (e.g., Best vs Acceptable), reflecting the models’ growing inability to distinguish between relevance grades on such higher difficulty tasks.
Trend 2: Indecision vs. Hallucination.
The most revealing commonality across all models is the source of error. By comparing “Acc” and “Acc (No Ties)” in Table 2, a consistent behavioral pattern emerges: LLMs are accurate when they express a decisive preference, while overrating primarily arises from their tendency to assign equivocal judgments rather than making clear distinctions. Put simply, when ties are excluded, all models exhibit strong reliability. Qwen achieves an impressive Acc (No Ties) of 93% on DL2020, and even the lowest performing model exceeds 85%. These findings suggest that LLMs do not typically hallucinate that a bad document is better than a good one (as demonstrated by the rather low “Wrong Wins” rate); rather, they fail to perceive the quality gap between the good and bad passages, defaulting to a Tie.
The confidence difference between correct and incorrect judgments is negligible in the pairwise scenario. On DL2019, Gemma’s average confidence drops only marginally from 0.979 on correct judgments (Corr) to 0.964 on incorrect ones (Incorr); Llama shows an even smaller gap. In other words, LLM confidence is virtually indistinguishable, regardless of judgment correctness.
@@ 3.1. Semantic-Preserving Structural Variations
LLMs are known to prefer longer responses regardless of quality (Zheng et al., 2023; Saito et al., 2023). We generate summarized (SUM) and expanded (EXP) passage variants while ensuring semantic equivalence. Comparing relevance scores across these variants reveals whether passage length is mistakenly treated as a relevance signal.
@@
Findings from Balog et al. (2025) indicate no general bias from LLM-judges to LLM-generated content, even within the same model family. We use Gemini-2.5-Flash to generate stylistic and length-based content variations. Unlike Balog et al. (2025), who assume rewriting preserves relevance (validated by human assessors (Dai et al., 2024)), we perform an explicit quality-check by prompting Gemini-2.5-Flash to verify that all rewritten passages preserve the meaning of the original passage. Across all prompts, we retain only the passages identified as “meaning preserved”.
@@
For each query, we independently sample 20 passages per predicted relevance label (0–3) for each LLM. For each passage, we generate four syntactic variants and assess label consistency across versions to analyze intrinsic model preferences. To avoid redundant rewriting from overlapping samples across LLMs, passages are pooled only at the rewriting stage, while each LLM is re-judged and evaluated solely on its own sampled query–rewritten passage pairs.
@@ Results.
Figure 1 shows the results as label transition matrices, capturing the change in judgment due to the four different passage re-writing treatments as compared to the original (unedited passage) LLM judgment. Across all models, passive voice rewrites and active voice rewrites produce nearly identical effects on LLM judgments, indicating no systematic preference for one form over the other. Summarization consistently reduces relevance label across models, whereas passage expansion systematically inflates relevance label for originally low-relevance passages – but can still degrade labels for highly relevant passages. These results indicate that even when semantic meaning is preserved, LLM-based relevance judgments exhibit a clear preference for longer passages.

### v3

@@ Abstract.
Human relevance assessment is time-consuming and cognitively intensive, limiting the scalability of Information Retrieval evaluation. This has led to growing interest in using large language models (LLMs) as proxies for human judges. However, it remains an open question whether LLM-based relevance judgments are reliable, stable, and rigorous enough to match humans for relevance assessment. In this work, we conduct a study of overrating behavior in LLM-based relevance judgments across model backbones, evaluation paradigms (pointwise and pairwise), and passage modification strategies. We show that models consistently assign inflated relevance scores — often with high confidence — to passages that do not genuinely satisfy the underlying information need, revealing a system-wide bias rather than random fluctuations in judgment. Furthermore, controlled experiments show that LLM-based relevance judgments can be highly sensitive to passage length and surface-level lexical cues. These results raise concerns about the usage of LLMs as drop-in replacements for human relevance assessors, and highlight the urgent need for careful diagnostic evaluation frameworks when applying LLMs for relevance assessments. Our code and results are publicly available. [^1] https://github.com/chutingyu/Exploring-Overrating
@@
journalyear: 2026copyright: ccconference: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 20–24, 2026; Melbourne, VIC, Australiabooktitle: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’26), July 20–24, 2026, Melbourne, VIC, Australiadoi: 10.1145/3805712.3809905isbn: 979-8-4007-2599-9/2026/07ccs: Information systems Relevance assessment
@@ 1. Introduction and Background
Building reliable relevance annotations for evaluation has long been a central challenge in Information Retrieval (IR). Cranfield-style collections (Cleverdon, 1967), such as TREC DL (Craswell et al., 2020, 2021), rely on trained assessors to annotate query-document pairs, costing substantial time and effort (Soboroff, 2025), while crowdsourced annotations often suffer from limited reliability (Guo et al., 2023; Rahmani et al., 2025; Damessie et al., 2017). Large language models (LLMs) offer a potential alternative by enabling relevance judgments to be generated rapidly and at scale.
Recent studies report that LLM-based relevance judgments can closely align with human assessments. Thomas et al. (2024) employed real searcher feedback to select LLM–prompt configurations aligned with user preferences, demonstrating performance comparable to human assessors in identifying strong systems and difficult queries. These findings were reproduced and extended by Upadhyay et al. (2024a), who further released the open-source UMBRELA framework (Upadhyay et al., 2024b).  Upadhyay et al. (2025) showed that LLM-generated judgments can replace traditional human labels at the run level in the TREC 2024 RAG Track, with limited benefit from additional human-in-the-loop intervention.
While LLMs have shown to be effective annotation tools, their potential shortcomings as relevance assessors remains insufficiently understood. LLM-based judgments raise several issues. For example, Alaofi et al. (2024) show that LLMs are vulnerable to keyword stuffing, systematically overrating non-relevant passages that contain query terms by assigning inflated relevance labels. Inflated relevance labels, as compared to humans, have been reported in prior work (Upadhyay et al., 2025; Meng et al., 2025). This behavior also demonstrates how strong aggregate performance can obscure judgment-level errors, which can directly impact downstream evaluation tasks. Moreover, using LLMs as both a retrieval method and as an evaluation (labeling) method introduces circularity, confounding true system improvements with LLM-specific preferences (Clarke and Dietz, 2024). These issues highlight the potential risks when adopting LLMs as relevance assessors.
Since IR systems are measured over the set of relevance labels for a given set of topics, these labels serve as an upper bound on measurable effectiveness. When relevance judgments are generated by LLMs, this upper bound is implicitly shaped by the models’ own judgment behaviors. LLM-based relevance assessments tend to exhibit label inflation (e.g., assigning a document a higher relevance label than that assigned by a human assessor), which may distort evaluation signals and mask genuine differences in retrieval quality. However, it remains unclear how widespread this behavior is in different model families and evaluation settings.
@@ 2. Overrating Behavior in LLM Judges
Beyond label inflation, we analyze the token-level confidence on the LLM’s predicted label or preference. In the pointwise setting, we determine confidently incorrect prediction as cases where the model assigns near-deterministic probabilities (i.e., softmax probabilities approaching 1.0) to incorrect label tokens, which are measured by the normalized logits on the generation of the label. In the pairwise setting, instability manifests where the LLM yields varying preferences under positional swapping (e.g., $A>B$ vs. $B>A$) while still assigning high confidence to the assessments.
@@ 2.1. Experimental Setup
We conduct experiments on the TREC Deep Learning (DL) Track passage ranking tasks from 2019-2020 (Craswell et al., 2020, 2021), using human relevance labels (0–3) as ground truth. We evaluate four open-weight LLMs: Llama-3.2-3B (Grattafiori et al., 2024), Gemma-3-4B (Team, 2025), Mistral-7B (Jiang et al., 2023), and Qwen-3-8B (Yang et al., 2025) under both pointwise and pairwise settings.
@@
Pointwise Evaluation.
For pointwise judgments, we measure label inflation through two metrics: (1) Overrate Ratio, the proportion of the instances where the LLM assigns a relevance label exceeding the human label; and (2) Mean Bias, the average signed difference between LLM-assigned labels and human relevance labels (LLM minus Human). We additionally report Cohen’s $\kappa$ to analyze agreement with human judgments. These metrics collectively quantify both the magnitude and distributional patterns of label inflation exhibited by different LLMs. To understand model confidence, we binarized both the human and LLM judgments according to the TREC standard. [^3] Not relevant: 0 and $1$; Relevant: $\geq 2$. See: https://trec.nist.gov/data/deep2019.html Then, we aggregate the token-level confidence according to the True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) classes when comparing the human and LLM judgment pairs. Our intuition is that models may be less confident for incorrect (FP or FN) judgments.
Pairwise Evaluation.
For pairwise judgment, passages are grouped per query into three categories: Best (top‐graded relevant), Acceptable (relevant, but not the highest grade), and Unacceptable (non‐relevant) (Arabzadeh and Clarke, 2025). We sampled 20 pairs per query from each of the following settings: Best vs. Acceptable; Acceptable vs. Unacceptable; and Best vs. Unacceptable. We treat the higher‐class passage as the ground‐truth winner, and pairs are judged in both presentation orders. We force the LLM to prefer one of the two documents, rather than allowing an explicit tie option. As such, we define a tie as the case where a contradiction arises between the two presentation orderings, meaning that the judgments cannot be linearized. This indicates that the model cannot reliably distinguish the relevance difference between the passages. Therefore, we interpret such cases as the model treating the two passages as equally relevant.
@@
2.2. Results: Pointwise Judgment
Model Overrated (%) Underrated (%) Mean Bias Cohen’s $\kappa$ Avg Confidence
Bin. Grad. Bin. Grad. Bin. Grad. Bin. Grad. TP FP TN FN
DL19 Gemma 37.81 61.84 3.74 8.36 0.341 0.736 0.246 0.107 0.987 0.989 0.989 0.977
Llama 18.82 49.72 11.09 12.17 0.077 0.564 0.304 0.163 0.902 0.881 0.958 0.914
Mistral 21.53 54.23 7.45 10.45 0.141 0.557 0.369 0.150 0.939 0.932 0.951 0.904
Qwen 20.62 47.02 6.10 10.3 0.145 0.459 0.421 0.215 0.977 0.976 0.991 0.986
DL20 Gemma 44.78 66.51 1.66 5.36 0.431 0.895 0.163 0.088 0.983 0.979 0.986 0.981
Llama 18.57 51.11 6.21 8.00 0.124 0.691 0.265 0.149 0.908 0.870 0.920 0.871
Mistral 22.32 55.08 4.35 7.47 0.180 0.646 0.293 0.140 0.941 0.925 0.937 0.910
Qwen 23.42 44.59 2.79 6.61 0.206 0.516 0.338 0.235 0.977 0.974 0.986 0.968
Table 1. Pointwise overrating and underrating metrics on DL2019 and DL2020 under binary and graded relevance labels. Average Confidence is computed under the graded relevance label setting.
Table 1 demonstrates the pervasive nature of overrating across all models and datasets from our results. We consistently observe that the overrating ratio is substantially higher than the underrating ratio, indicating LLMs rarely assign labels lower than human annotators. Since the two settings use different relevance scales (Section 2.1), we binarized graded labels based on TREC standard to enable a fair comparison; the observed overrating ratio becomes lower than that obtained from direct binary evaluation. This suggests that a substantial portion of overrating errors in the graded setting occurs within the relevant or highly relevant categories. Cohen’s $\kappa$ is consistently lower in the graded setting than in the binary setting, which suggests that LLM judgments are less consistent with human annotations in finer-grained relevance, although agreement is typically lower in graded scenarios due to the additional freedom afforded by having more relevance levels.
Examining the label transition matrices in the graded setting also reveals that the overrating behavior is dominated by soft errors: non-relevant passages are rarely assigned the highest relevance score, but are frequently misclassified as marginally or moderately relevant. This systematic reluctance to assign the non-relevant label inflates relevance estimates under graded evaluation.
@@ 2.3. Results: Pairwise Judgment
Trend 1: Lack of Discrimination.
As shown in Table 2, the “Tie Rate” is substantial for all models, ranging from approximately 24% to over 45%. This indicates order-induced inconsistency (preference flips under swapping), not consistent errors where the model repeats the same wrong choice. Furthermore, we investigated how judgment varies with the different graded pairs, as, intuitively, the Best vs Unacceptable pairs should be the easiest to discriminate. However, we observed that all four LLMs tested consistently overrate “Unacceptable” passages, treating them as qualitatively equivalent to the “Best” passages in nearly one-third of cases. Moreover, the tie rate increases systematically as the discrimination task becomes more difficult (e.g., Best vs Acceptable), reflecting the models’ growing inability to distinguish between relevance grades on such higher difficulty tasks.
Trend 2: Indecision vs. Hallucination.
The most revealing commonality across all models is the source of error. By comparing accuracy and accuracy (no ties) in Table 2, a consistent behavioral pattern emerges: LLMs are accurate when they express a decisive preference, while overrating primarily arises from their tendency to assign equivocal judgments rather than making clear distinctions. Put simply, when ties are excluded, all models exhibit strong reliability. Qwen achieves an impressive accuracy (no ties) of 93% on DL2020, and even the lowest performing model exceeds 85%. These findings suggest that LLMs do not typically hallucinate that a bad passage is better than a good one (as demonstrated by the rather low “Wrong Wins” rate); rather, they fail to perceive the quality gap between the good and bad passages, defaulting to a Tie.
The confidence difference between correct and incorrect judgments is negligible in the pairwise scenario. On DL2019, Gemma’s average confidence drops only marginally from 0.979 on correct judgments (Corr.) to 0.964 on incorrect ones (Incorr.); Llama shows an even smaller gap. In other words, LLM confidence is virtually indistinguishable, regardless of judgment correctness.
@@
Acc. Acc. (No Ties) Tie Rate Wrong Wins Corr. Incorr.
@@
Table 2. Pairwise preference-judgment results on DL2019 and DL2020. Average confidence is computed on no-tie cases only for correct and incorrect preferences.
@@ 3.1. Semantic-Preserving Structural Variations
LLMs are known to prefer longer responses regardless of quality (Zheng et al., 2023; Saito et al., 2023), and input length shifts LLM label distributions in model-dependent ways (Mohtadi et al., 2026). We generate summarized (SUM) and expanded (EXP) passage variants while ensuring semantic equivalence. Comparing relevance scores across these variants reveals whether passage length is mistakenly treated as a relevance signal.
@@
Findings from Balog et al. (2025) indicate no general bias from LLM-judges to LLM-generated content, even within the same model family. We use Gemini-2.5-Flash to generate stylistic and length-based content variations. Unlike Balog et al. (2025), who assume rewriting preserves relevance (validated by human assessors (Dai et al., 2024)), we perform an explicit quality-check by prompting Gemini-2.5-Flash to verify that all rewritten passages preserve the meaning of the original passage. Across all prompts, we retain only the passages identified as “meaning preserved.”
@@
For each query, we independently sample (at maximum) 20 passages per predicted relevance label (0–3) for each LLM. For each passage, we generate four syntactic variants and assess label consistency across versions to analyze intrinsic model preferences. To avoid redundant rewriting from overlapping samples across LLMs, passages are pooled only at the rewriting stage, while each LLM is re-judged and evaluated solely on its own sampled query–rewritten passage pairs.
@@ Results.
Figure 1 shows the results as label transition matrices, capturing the change in judgment due to the four different passage re-writing treatments as compared to the original (unedited passage) LLM judgment. Across all models, passive voice rewrites and active voice rewrites produce nearly identical effects on LLM judgments, indicating no systematic preference for one form over the other. Summarization consistently reduces relevance labels across models, whereas passage expansion does not consistently inflate relevance labels. Passage expansion systematically inflates relevance labels for originally low-relevance passages – but can still degrade labels for highly relevant passages. These results indicate that even when semantic meaning is preserved, LLM-based relevance judgments exhibit a clear tendency to disfavor short passages.